Y=2x² - 12x +21
Axis of symmetry X=
Vertex ( , )
1
Expert's answer
2019-10-01T09:59:29-0400

If the parabola is given as y = a x ^ 2 + b x + c,

then the axis of symmetry is located as a straight line perpendicular to the X axis and passing through the vertex.

To find the abscissa of the vertex, we need to find the derivative: y '= 2ax + b and set it to zero, whence we get x = -b / 2a, x = 12/4 = 3; x = 3 is the axis of symmetry.